Circular No. 8340

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004br IN NGC 4493 Further to IAUC 8336, J. Graham and R. J. Foley report the LOSS

discovery, on KAIT images taken on May 15.3 and 16.3 UT, of an apparent supernova (mag ∼ 16.3) located at α = 12^{h}31^{m}07^{s}.80, δ = +0^{o}36′47′′.6 (equinox 2000.0), which is 9′′.0 west and 0′′.6 south of the nucleus of NGC 4493. A KAIT image taken on Apr. 29.3 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2003 Q7–Q9, R2–R5, S5-S9, T8, U4 (SOHO) Further to IAUC 8339, K. Battams reports measurements for addi-

tional near-sun comets found on SOHO website images by M. Boschat (C/2003 Q7, Q8, S6), J. Sachs (C/2003 Q9, R3, S7, S8), X.-m. Zhou (C/2003 R2), R. Kracht (C/2003 R4, S9), K. Černis (C/2003 R5), X. Leprette (C/2003 S5), and R. Matson (C/2003 T8, U4). C/2003 Q8, Q9, R2, R3, S5, S6, and S7 were observed only in C3 images, while C/2003 R4, S9, T8, and U4 were observed only in C2 images; the remaining comets were observed in both coronagraphs. The reduced astrometry by B. G. Marsden and R. Kisala, and orbital elements by Marsden, appear on the MPECs cited below. C/2003 U4 belongs to the Meyer group, and C/2003 R4 and S9 belong to no known group. C/2003 R5 belongs to a new group that certainly includes C/1999 R1 and probably includes C/2002 R5. The remaining objects are Kreutz sungrazers.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 Q7 Aug. 28.988  9^{h}57.^{m}7  + 8^{o}14′  2004-J57
C/2003 Q8  29.388  10 10.8  + 8 46  2004-J57
C/2003 Q9  30.613  10 10.3  + 7 52  2004-J57
C/2003 R2 Sept. 5.388  10 38.5  + 6 24  2004-J57
C/2003 R3  8.763  10 44.9  + 4 51  2004-J57
C/2003 R4  7.021  11 01.9  + 8 01  2004-J59
C/2003 R5  8.204  11 01.2  + 6 07  2004-J59
C/2003 S5  18.388  11 20.5  + 1 12  2004-J58
C/2003 S6  18.571  11 26.5  + 1 21  2004-J58
C/2003 S7  21.363  11 34.2  + 0 03  2004-J58
C/2003 S8  25.763  11 46.7  − 1 57  2004-J58
C/2003 S9  25.021  12 01.1  + 1 08  2004-J59
C/2003 T8 Oct. 8.963  12 48.1  − 6 37  2004-J58
C/2003 U4  21.079  13 37.3  − 8 51  2004-J21
2004 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green